OpenStudy (anonymous):

What is the lateral area of a rectangular prism whose length is 2 cm, width is 3 cm, and height is 6 cm?

OpenStudy (auctoratrox):

the lateral area is the sum of the areas of all the faces that make up the prism. the base area is 2x3 = 6 and all the triangles have areas of (1/2)base x height, so we get 6 + 6 + 6 + 9 + 9 = 36 squared centimeters.

OpenStudy (auctoratrox):

oh wait, it's not a pyramid, my mistake. so the lateral area is the sum of all the sides, so we get 6x3x2 plus 6x2x4 = 84 squared centimeters.

Directrix (directrix):

The *lateral area* of a rectangular prism does not include the two bases of the prism. The lateral area will be the sum of the four rectangular lateral faces. LA of Prism = the perimeter of the base times the height. LA = (2 + 3 + 2 + 3) *6 = 10*6 = 60 square cm